# CSC-MPPI: A Novel Constrained MPPI Framework with DBSCAN for Reliable Obstacle Avoidance*

### v1

@@ Abstract
Figure 1: Comparison of standard MPPI and CSC-MPPI. (a) Standard MPPI results, where the sampled trajectories (red solid lines) are predominantly located in high-cost regions, resulting in a suboptimal path (cyan solid line) that leads to collision with the obstacle. (b) CSC-MPPI first adjusts the particles to ensure they satisfy constraints and then clusters the trajectories (red and orange solid lines) based on spatial proximity and cost similarity. Trajectories that do not belong to any cluster are considered noise trajectories (gray solid lines). (c) An optimal trajectory is selected from each cluster (light blue and yellow dashed lines). (d) Finally, the trajectory with the lowest cost among clusters is selected as the final optimal trajectory (cyan solid line).
@@ I-A Related Works
To address this problem, Yan and Devasia [10] introduced Output-Sampled MPPI (o-MPPI), which improves constraint satisfaction by selectively sampling within predefined acceptable regions in the output space. However, o-MPPI does not provide a clear strategy for handling samples that fall outside these acceptable regions. Balci et al. [11] proposed Constrained Covariance Steering-Based MPPI (CCS-MPPI), which enforces state and input constraints through hyperplane-based method. Nonetheless, its reliance on probabilistic constraints can lead to occasional violations. To strengthen constraint handling in MPPI, Yin et al. [12] developed Shield-MPPI, which combines a dual-layer safety mechanism with Control Barrier Functions (CBFs), a cost function that penalizes unsafe trajectories, and a reactive safety shield that adjusts controls in real time. Although this structured approach enhances safety, it still depends on soft constraints and cannot fully guarantee constraint satisfaction. Finally, Tao et al. [13] introduced SCBF-MPPI, which embeds stochastic control barrier functions (SCBFs) directly into the sampling process, dynamically shifting the mean and covariance of the sampling distribution to ensure trajectories remain within safe regions. However, SCBF-MPPI also operates probabilistically, thus absolute constraint enforcement remains unattainable.
@@ I-B Overview of Our Approach
To overcome the limitations of conventional MPPI in constraint enforcement, we propose Constrained Sampling Cluster MPPI (CSC-MPPI), which integrates explicit hard constraint satisfaction into the MPPI framework while preserving its adaptability. The proposed approach consists of a sequential process that refines the standard MPPI sampling strategy to ensure feasibility.
@@
Therefore, the main advantages of the proposed control framework are as follows: First, unlike conventional MPPI methods that rely on soft constraints, CSC-MPPI explicitly enforces hard constraints on both states and control inputs, ensuring feasibility in highly constrained environments. To the best of the authors’ knowledge, this work represents the first attempt to impose hard constraints on both state and control inputs within the MPPI framework. Second, the integration of DBSCAN eliminates the risk of constraint violations caused by weighted averaging. In addition, DBSCAN prevents sample clustering in high-cost regions, reducing the risk of local minima and improving trajectory efficiency. Finally, extensive simulation and real-world experiments on mobile robot obstacle avoidance demonstrate that CSC-MPPI robustly handles constraints and consistently achieves safe and efficient trajectories.
Overall, the experimental results demonstrate that CSC-MPPI reliably enforces hard constraints and optimizes trajectories in challenging obstacle avoidance scenarios. These findings underscore its potential for real-time applications across diverse robotic platforms.
The remainder of this paper is as follows. Section II reviews the MPPI algorithm and its problem formulation. Next, we present the formulation of CSC-MPPI with primal-dual gradient method and DBSCAN in Sec. III. Section IV describes the experimental validations of the proposed method. Finally, the paper is concluded in Sec. V.
@@ II-B Review of the MPPI
${\eta}=\sum_{k=0}^{K-1}\exp\left(-\frac{1}{\lambda}S(\prescript{k}{}{\boldsymbol{V}})-\sum_{t=0}^{N-1}(\hat{\boldsymbol{u}}_{t}-\tilde{\boldsymbol{u}}_{t})^{T}\boldsymbol{\Sigma}_{\boldsymbol{u}}^{-1}\prescript{k}{}{\boldsymbol{v}}_{t}\right)d\boldsymbol{V},$ (8)
@@ III CONSTRAINED SAMPLING CLUSTER MPPI
As discussed in Section I, MPPI often struggles to handle hard constraints on states and control inputs. Since trajectory updates in MPPI emerge from stochastic sampling and weighted averaging, naive application can lead to locally suboptimal or even infeasible solutions—particularly when sampled trajectories cluster in high-cost regions. In such scenarios, MPPI may become trapped in local minima, generate suboptimal paths, or increase the risk of collisions. To address these limitations, this paper introduces CSC-MPPI, which integrates a primal-dual gradient-based adjustment with DBSCAN clustering. As illustrated in Fig. 2, the approach begins by drawing trajectory samples from a proposal distribution, similar to standard MPPI. Instead of keeping these trajectories unconstrained, a primal-dual gradient method iteratively refines them to ensure compliance with prescribed state and control bounds. Once feasibility is ensured, DBSCAN clusters the resulting control candidates based on spatial proximity and cost similarity. Clustering serves to mitigate the risk that arises when the weighted averaging process in MPPI undermines the imposed constraints. Although the sampling process ensures constraint satisfaction initially, weighted averaging over samples with similar costs may degrade constraint adherence. By forming clusters that account for both cost similarity and spatial proximity, DBSCAN reduces this risk and maintains feasibility. From each cluster, the control input with the lowest cost, as determined by the MPPI cost function, is selected for execution.
@@ III-A Primal-Dual Gradient-Based Constraint Enforcement
As shown in Fig. 2c, the main idea of the proposed method is to shift the randomly sampled candidates into the feasible region defined by the prescribed constraints. Formally, this can be expressed as follows:
@@
$s.t.\quad\underline{\boldsymbol{v}}\leq\prescript{k}{}{\boldsymbol{v}}_{t}\leq\overline{\boldsymbol{v}},$
@@
${\nabla}_{\prescript{k}{}{\boldsymbol{v}}_{t}}{g}(\prescript{k}{}{\boldsymbol{x}}_{t+1})=\frac{\partial{g}(\prescript{k}{}{\boldsymbol{x}}_{t+1})}{\partial{f}(\prescript{k}{}{\boldsymbol{x}}_{t},\prescript{k}{}{\boldsymbol{v}}_{t})}\frac{\partial{f}(\prescript{k}{}{\boldsymbol{x}}_{t},\prescript{k}{}{\boldsymbol{v}}_{t})}{\partial\prescript{k}{}{\boldsymbol{v}}_{t}}$ (12)
@@
$\prescript{k}{}{\boldsymbol{v}}^{\text{new}}_{t}=\prescript{k}{}{\boldsymbol{v}}_{t}-\boldsymbol{\alpha}\circ{\nabla}_{\prescript{k}{}{\boldsymbol{v}}_{t}}\prescript{k}{}{\mathcal{L}}_{t}(\prescript{k}{}{\boldsymbol{v}_{t}},\prescript{k}{}{\underline{\boldsymbol{\mu}}}_{t},\prescript{k}{}{\overline{\boldsymbol{\mu}}}_{t})$ (14)
@@ III-B Clustering Sampled Trajectories Using DBSCAN
Algorithm 1 summarizes the overall flow of the CSC-MPPI framework, detailing the sequential steps from initial trajectory sampling and constraint adjustment via the primal-dual gradient method to subsequent clustering using DBSCAN and the final selection of the optimal control input.
@@
18:              $\prescript{k}{}{\boldsymbol{v}}_{t}\leftarrow\prescript{k}{}{\boldsymbol{v}}_{t}+\boldsymbol{\alpha}\circ{\nabla}_{\prescript{k}{}{\boldsymbol{v}_{t}}}\prescript{k}{}{\mathcal{L}}_{t}(\prescript{k}{}{\boldsymbol{v}_{t}},\prescript{k}{}{\underline{\boldsymbol{\mu}}}_{t},\prescript{k}{}{\overline{\boldsymbol{\mu}}}_{t})$
@@ IV EXPERIMENTS
In this section, we validate the effectiveness of the proposed method through simulations and real-world experiments. In the simulation environment, we compare the proposed method with standard MPPI in obstacle avoidance task. Additionally, to examine the risk of constraint violation due to weighted averaging, we evaluate CSC-MPPI with and without DBSCAN. Finally, the real-world experiments were conducted to validate CSC-MPPI in practical scenarios.
IV-A Simulation Setup
@@
Figure 3: Simulation environments for performance comparison. In both scenarios, the purple circle represents the goal position, and the black circle indicates the initial position of the robot. Gray circles represent obstacles, with dashed gray circles denoting the boundaries that account for the robot’s footprint. (a) Environment #1, which includes one dynamic and two static obstacles, designed to compare the performance of standard MPPI and CSC-MPPI. (b) Environment #2, a simplified environment with only one static obstacle, used to evaluate the effect of DBSCAN in CSC-MPPI.
$\begin{split}x_{t+1}&=x_{t}+u_{v_{t}}\cos\theta_{t}\,dt\\ y_{t+1}&=y_{t}+u_{v_{t}}\sin\theta_{t}\,dt\\ \theta_{t+1}&=\theta_{t}+u_{\omega_{t}}\,dt\end{split}$ (15)
The system state $\boldsymbol{x}=[x,y,\theta]^{T}$ represents the position and orientation of the robot, while the control input $\boldsymbol{u}=[u_{v},u_{w}]^{T}$ represents the linear and angular velocity commands. The simulations were conducted on a computer equipped with an Intel Core i7-13700F processor running at 2.1 GHz, 32 GB of RAM, and an NVIDIA GeForce RTX 4060 GPU. Parallel computation was performed using a Python-based simulation framework. For simulation environment #1, the temperature parameter $\lambda$ was set to 0.01, while for simulation environment #2, $\lambda$ was set to 0.7. The remaining MPPI parameters were kept identical across both environments. The covariance matrix $\boldsymbol{\Sigma}_{\boldsymbol{u}}$ was defined as $\text{diag}(\sigma_{u_{v}}^{2},\sigma_{u_{w}}^{2})=\text{diag}(0.1^{2},1.0^{2})$. The time horizon $N$ was set to 30 with a time step of $dt=0.03$. The linear velocity was constrained between 00 m/s and $0.5$ m/s, while the angular velocity was constrained between $-3.0$ rad/s and $3.0$ rad/s. The tolerance for the goal state is defined as $\tau_{p}=0.1$ for the position $(x,y)$ and $\tau_{\theta}=0.2$ for the orientation $\theta$. The running cost $l$ is defined as
@@
IV-B Simulation Scenarios
As shown in Fig. 3a, Environment #1 is designed to compare the performance of standard MPPI and CSC-MPPI. The autonomous robot starts from the initial state $\boldsymbol{x}_{s}=[-1.0,-1.0,\pi/2]^{T}$ and must navigate towards the goal state $\boldsymbol{x}_{f}=[2.0,2.0,\pi/2]^{T}$ while avoiding obstacles. The environment includes one dynamic obstacle and two static obstacles. The dynamic obstacle has a radius of $0.3$ m and moves from $[-1.0,0.0]^{T}$ to $[0.5,0.0]^{T}$ with a constant velocity of $0.53$ m/s. The static obstacles are centered at $[0.0,1.0]^{T}$ and $[1.5,0.7]^{T}$, with radii of $0.4$ m and $0.5$ m, respectively.
As shown in Fig. 3b, Environment #2 is designed to evaluate the performance of CSC-MPPI with and without DBSCAN. The robot’s initial state is $\boldsymbol{x}_{s}=[-1.0,0.0,0.0]^{T}$, and the goal state is $\boldsymbol{x}_{f}=[1.0,0.0,0.0]^{T}$. The environment consists of a single static obstacle centered at $[0.0,0.0]^{T}$ with a radius of 0.5 m. The number of sampled trajectories is set to $K=300$.
@@ IV-C Simulation Results
The computational efficiency of each method was also analyzed. Despite increasing the number of sampled trajectories from 20 to 300, both standard MPPI and CSC-MPPI maintained relatively stable computation times ranging from 1.669 ms to 1.739 ms for standard MPPI and from 8.870 ms to 9.007 ms for CSC-MPPI. This consistent performance is attributed to GPU-based parallel computation, which mitigates the computational burden even as sample size increases. The additional computational overhead in CSC-MPPI was primarily due to the gradient method. Clustering, required only an average of 6.03 ms for $K=20$, 6.06 ms for $K=50$, and 5.99 ms for $K=300$, a duration comparable to the total computation time of standard MPPI. Despite the increased computational cost, CSC-MPPI exhibited superior performance in collision avoidance and trajectory efficiency while utilizing only 10% of the samples required by standard MPPI.
Figure 5: Trajectory comparison of CSC-MPPI with and without DBSCAN in Environment #2. (a) The trajectories generated by CSC-MPPI without DBSCAN. The red lines and cyan line represent sampled trajectories and the optimal trajectory obtained from (6), respectively. (b) The trajectories generated by CSC-MPPI with DBSCAN. The red, yellow, orange, and green lines represent clustered trajectories. The cyan line indicates the final optimal trajectory.
In Environment #2, the experiment was conducted to evaluate constraint satisfaction in CSC-MPPI with and without DBSCAN. Both approaches achieved a 0% collision rate. However, in terms of constraint satisfaction for the trajectory corresponding to the optimal input sequence obtained from (6), the satisfaction rate was 100% with DBSCAN, whereas it dropped to 80% without DBSCAN. As shown in Fig. 5a, even though the sampled trajectories satisfy the constraints, if multiple sampled trajectories have similar costs, the weighted averaging process may result in a trajectory that does not satisfy the constraints. In contrast, as illustrated in Fig. 5b, CSC-MPPI with DBSCAN selects a single cluster from multiple groups, ensuring that the final optimal input sequence satisfies the constraints. Thus, the proposed method effectively eliminates the risk of constraint violation caused by weighted averaging.

### v2

@@ Abstract
Figure 1: Comparison of standard MPPI and CSC-MPPI. (a) Sampled trajectories from standard MPPI (red solid lines) are concentrated in high-cost regions, producing a suboptimal path (cyan solid line) that results in collision. (b) CSC-MPPI adjusts samples and clusters feasible trajectories (red/orange solid lines), while noise samples are shown in gray. (c) A representative trajectory (light blue/yellow dashed lines) is selected from each cluster. (d) The final optimal trajectory (cyan solid line) is chosen as the lowest-cost candidate.
@@ I-A Related Works
To address this problem, Yan and Devasia [10] introduced Output-Sampled MPPI (o-MPPI), which improves constraint satisfaction by selectively sampling within predefined acceptable regions in the output space. However, o-MPPI does not provide a clear strategy for handling samples that fall outside these acceptable regions. Balci et al. [11] proposed Constrained Covariance Steering-Based MPPI (CCS-MPPI), which enforces state and input constraints through hyperplane-based method. Nonetheless, its reliance on probabilistic constraints can lead to occasional violations. To strengthen constraint handling in MPPI, Yin et al. [12] developed Shield-MPPI, which combines a dual-layer safety mechanism with Control Barrier Functions (CBFs), a cost function that penalizes unsafe trajectories, and a reactive safety shield that adjusts controls in real time. Although this structured approach enhances safety, it still depends on soft constraints and cannot fully guarantee constraint satisfaction. Finally, Tao et al. [13] introduced Stochastic Control Barrier Function MPPI (SCBF-MPPI), which embeds Stochastic Control Barrier Functions (SCBFs) directly into the sampling process, dynamically shifting the mean and covariance of the sampling distribution to ensure trajectories remain within safe regions. However, SCBF-MPPI also operates probabilistically, thus absolute constraint enforcement remains unattainable.
@@ I-B Overview of Our Approach
To overcome the limitations of conventional MPPI in constraint enforcement, we propose CSC-MPPI, which integrates explicit hard constraint satisfaction into the MPPI framework while preserving its adaptability. The proposed approach refines the standard MPPI sampling strategy through a sequential process to ensure feasibility under strict state and input constraints.
@@
The main contributions of this paper are as follows:
  • Hard constraint enforcement: Unlike conventional MPPI methods that rely on soft penalties, CSC-MPPI explicitly enforces hard constraints on both state and control inputs. To the best of the authors’ knowledge, this is the first MPPI-based approach to guarantee hard constraint satisfaction in both domains.
  • Robust control input selection via clustering: By employing DBSCAN, CSC-MPPI prevents high-cost or noisy trajectories from affecting control selection. This addresses the critical limitation of weighted averaging in standard MPPI, which often leads to constraint violations.
  • Improved safety and efficiency: The combination of constraint-aware sampling and clustering leads to trajectories that are not only feasible but also more efficient, reducing the risk of local minima. This has been consistently validated in both simulation and real-world navigation experiments.
The remainder of this paper is organized as follows. Section II reviews the MPPI algorithm and formulates the control problem. Section III presents the proposed CSC-MPPI method in detail. Section IV describes experimental validations through simulations and hardware tests. Finally, conclusions are drawn in Section V.
@@ II-B Review of the MPPI
${\eta}=\sum_{k=0}^{K-1}\exp\left(-\frac{1}{\lambda}S(\prescript{k}{}{\boldsymbol{V}})-\sum_{t=0}^{N-1}(\hat{\boldsymbol{u}}_{t}-\tilde{\boldsymbol{u}}_{t})^{T}\boldsymbol{\Sigma}_{\boldsymbol{u}}^{-1}\prescript{k}{}{\boldsymbol{v}}_{t}\right),$ (8)
@@ III CONSTRAINED SAMPLING CLUSTER MPPI
As discussed in Section I, MPPI often struggles to handle hard constraints on states and control inputs. Since trajectory updates in MPPI emerge from stochastic sampling and weighted averaging, naive application can lead to locally suboptimal or even infeasible solutions—particularly when sampled trajectories cluster in high-cost regions. In such scenarios, MPPI may become trapped in local minima, generate suboptimal paths, or increase the risk of collisions. To address these issues, the proposed CSC-MPPI framework combines a primal-dual gradient-based adjustment with DBSCAN clustering.
@@ III-A Primal-Dual Gradient-Based Constraint Enforcement
As shown in Fig. 2b and 2c, the main idea of the proposed method is to shift the randomly sampled candidates into the feasible region defined by the prescribed constraints. Formally, this can be expressed as follows:
@@
[…]
@@
${\nabla}_{\prescript{k}{}{\boldsymbol{v}}_{t}}{g}(\prescript{k}{}{\boldsymbol{x}}_{t+1})=\frac{\partial{g}(\prescript{k}{}{\boldsymbol{x}}_{t+1})}{\partial{f}(\prescript{k}{}{\boldsymbol{x}}_{t},\prescript{k}{}{\boldsymbol{v}}_{t})}\frac{\partial{f}(\prescript{k}{}{\boldsymbol{x}}_{t},\prescript{k}{}{\boldsymbol{v}}_{t})}{\partial\prescript{k}{}{\boldsymbol{v}}_{t}}.$ (12)
@@
$\prescript{k}{}{\boldsymbol{v}}^{\text{new}}_{t}=\prescript{k}{}{\boldsymbol{v}}_{t}-\boldsymbol{\alpha}\circ{\nabla}_{\prescript{k}{}{\boldsymbol{v}}_{t}}\prescript{k}{}{\mathcal{L}}_{t}(\prescript{k}{}{\boldsymbol{v}_{t}},\prescript{k}{}{\underline{\boldsymbol{\mu}}}_{t},\prescript{k}{}{\overline{\boldsymbol{\mu}}}_{t}),$ (14)
@@ III-B Clustering Sampled Trajectories Using DBSCAN
Algorithm 1 summarizes the overall CSC-MPPI pipeline. The initial sampled trajectories are generated via standard MPPI and propagated using the system dynamics (Lines 2–7). A primal-dual gradient update is then applied to each sample to satisfy state and control constraints (Lines 8–15). After computing the total cost of each trajectory (Lines 16–20), the control deviations and corresponding costs are stored as a dataset $\mathcal{D}=\{(\delta\boldsymbol{U}^{k},S^{k})\}$, which serves as input to DBSCAN (Lines 22-23). This clustering step groups the samples based on control similarity and cost proximity, while automatically discarding outliers. A softmax-weighted MPPI update is performed on each cluster (Lines 24–26), using only the samples within that cluster. Here, $\rho$ denotes the minimum cost within the cluster and serves as a reference to stabilize the weight computation. Finally, the cluster with the lowest-cost result is selected to generate the control output (Line 27).
@@
18:              $\prescript{k}{}{\boldsymbol{v}}_{t}\leftarrow\prescript{k}{}{\boldsymbol{v}}_{t}-\boldsymbol{\alpha}\circ{\nabla}_{\prescript{k}{}{\boldsymbol{v}_{t}}}\prescript{k}{}{\mathcal{L}}_{t}(\prescript{k}{}{\boldsymbol{v}_{t}},\prescript{k}{}{\underline{\boldsymbol{\mu}}}_{t},\prescript{k}{}{\overline{\boldsymbol{\mu}}}_{t})$
@@ IV EXPERIMENTS
In this section, we validate the effectiveness of the proposed method through simulations and real-world experiments. In the simulation environment, we compare the proposed method with standard MPPI in an obstacle avoidance task. Additionally, to examine the risk of constraint violation due to weighted averaging, we evaluate the impact of DBSCAN-based clustering by comparing a full version of the proposed CSC-MPPI and a simplified variant that omits the DBSCAN clustering step. Finally, the real-world experiments were conducted to validate CSC-MPPI in practical scenarios.
IV-A Simulation Scenarios
In both simulation environments shown in Fig. 3, the robot starts from the initial state indicated by a black circle and navigates toward the goal state marked by a purple circle. Obstacles are represented by gray circles, and the dashed boundaries around them reflect safety margins that account for the robot’s footprint.
As shown in Fig. 3a, Environment 1 is designed to compare the performance of standard MPPI and CSC-MPPI. The autonomous robot starts from $\boldsymbol{x}_{s}=[-1.0,-1.0,\pi/2]^{T}$ and must reach $\boldsymbol{x}_{f}=[2.0,2.0,\pi/2]^{T}$ while avoiding one dynamic and two static obstacles. The dynamic obstacle has a radius of 0.3 m and moves from $[-1.0,0.0]^{T}$ to $[0.5,0.0]^{T}$ at a constant velocity of 0.53 m/s. The static obstacles are centered at $[0.0,1.0]^{T}$ and $[1.5,0.7]^{T}$, with radii of 0.4 m and 0.5 m, respectively.
As shown in Fig. 3b, Environment 2 is a simpler scenario used to analyze the performance of CSC-MPPI and its variant without DBSCAN. The robot begins at $\boldsymbol{x}_{s}=[-1.0,0.0,0.0]^{T}$ and aims to reach $\boldsymbol{x}_{f}=[1.0,0.0,0.0]^{T}$, navigating around a single static obstacle centered at $[0.0,0.0]^{T}$ with a radius of 0.5 m.
IV-B Simulation Setup
@@
Figure 3: Simulation environments used for performance evaluation. (a) Environment #1 for comparison between standard MPPI and CSC-MPPI. (b) Environment #2 for analyzing the impact of DBSCAN.
$\begin{split}x_{t+1}&=x_{t}+u_{v_{t}}\cos\theta_{t}\,dt\\ y_{t+1}&=y_{t}+u_{v_{t}}\sin\theta_{t}\,dt\\ \theta_{t+1}&=\theta_{t}+u_{\omega_{t}}\,dt.\end{split}$ (15)
The system state $\boldsymbol{x}=[x,y,\theta]^{T}$ represents the position and orientation of the robot, while the control input $\boldsymbol{u}=[u_{v},u_{w}]^{T}$ represents the linear and angular velocity commands. The simulations were conducted on a computer equipped with an Intel Core i7-13700F processor running at 2.1 GHz, 32 GB of RAM, and an NVIDIA GeForce RTX 4060 GPU. Parallel computation was performed using a Python-based simulation framework. For simulation Environment #1, the temperature parameter $\lambda$ was set to 0.01, while for Environment #2, $\lambda$ was set to 0.7. This higher temperature in Environment #2 was deliberately chosen to highlight the effect of DBSCAN within the CSC-MPPI framework. In Environment #2, the number of sampled trajectories was fixed at $K=300$. The remaining MPPI parameters were kept identical across both environments. The covariance matrix $\boldsymbol{\Sigma}_{\boldsymbol{u}}$ was defined as $\text{diag}(\sigma_{u_{v}}^{2},\sigma_{u_{w}}^{2})=\text{diag}(0.1^{2},1.0^{2})$. The time horizon $N$ was set to 30 with a time step of $dt=0.03$. The linear velocity was constrained between 00 m/s and $0.5$ m/s, while the angular velocity was constrained between $-3.0$ rad/s and $3.0$ rad/s. The tolerance for the goal state is defined as $\tau_{p}=0.15$ m for the position $(x,y)$ and $\tau_{\theta}=0.25$ rad for the orientation $\theta$. The running cost $l$ is defined as
@@ IV-C Simulation Results
The computational efficiency of each method was also analyzed. Despite increasing the number of sampled trajectories from 20 to 300, both standard MPPI and CSC-MPPI maintained stable computation times ranging from 1.669 ms to 1.739 ms for standard MPPI and from 8.870 ms to 9.007 ms for CSC-MPPI. This consistent performance is attributed to GPU-based parallel computation, which mitigates the computational burden even as sample size increases. The increased computational time in CSC-MPPI is mainly caused by the primal-dual gradient adjustment and the DBSCAN-based clustering process. Specifically, performing DBSCAN clustering and obtaining representative sequences from the resulting clusters took an average of 6.03 ms for $K=20$, 6.06 ms for $K=50$, and 5.99 ms for $K=300$. Notably, CSC-MPPI achieves superior collision avoidance and trajectory efficiency while utilizing only 10% of the samples required by standard MPPI.
Figure 5: Trajectory comparison between CSC-MPPI and its variant without DBSCAN in Environment #2. (a) The trajectories generated by the CSC-MPPI variant without DBSCAN. The red lines and the cyan line represent the sampled trajectories and the optimal trajectory obtained from (6), respectively. (b) The trajectories generated by CSC-MPPI. Colored lines (e.g., red, yellow, orange, pink) represent clustered trajectories, and the cyan line indicates the final optimal trajectory.
In Environment #2, the experiment was conducted to evaluate constraint satisfaction by comparing CSC-MPPI and its variant without DBSCAN. Both approaches achieved a 0% collision rate. However, when considering constraint satisfaction of the trajectory corresponding to the optimal input sequence obtained from (6), the satisfaction rate was 100% for CSC-MPPI, whereas it dropped to 80% for the variant without DBSCAN. As shown in Fig. 5a, even though the sampled trajectories satisfy the constraints, if multiple sampled trajectories have similar costs, the weighted averaging process may result in a trajectory that does not satisfy the constraints. In contrast, as illustrated in Fig. 5b, CSC-MPPI uses DBSCAN to select a representative cluster from the set of feasible trajectories, thereby ensuring that the final optimal input sequence satisfies the constraints. These results demonstrate that the proposed method effectively mitigates the risk of constraint violation introduced by weighted averaging.